OpenStudy (anonymous):

1 What is the gravitational field strength at a height h above the surface of the Earth? R is the radius of the earth gR2(R+h)2 gR(R+h) Rg(R−h) R2g(R−h)2 2 The instrument used for experimental determination of the value of the universal gravitational constant G is called Galilio’s telescope ballistic pendulum Newton’s balance Cavendish balance 3 At what altitude above the earth's surface would the acceleration due to gravity be 49ms−2? Assume the mean radius of the earth is 64106 metres and the acceleration due to gravity 98ms−2 on the surface of the earth 26106m 323106m 465106m 776106m

OpenStudy (anonymous):

1. GM/(R+h)^2 That is nothing but \[(gR ^{2}) /( R+h)^{2} \] since \[g = GM/R ^{2} \] 3. acceleration due to gravity can never go to 49 ms^-2 above the earth's surface. It would go to 4.9 ms^-2 when R2/(R+h)2 = 1/2 determine h from this equation. R = radius of earth

OpenStudy (anonymous):

Thanks, so much

OpenStudy (anonymous):

which of the following has the same quantity as work. 1. momentum, 2. impulse, 3. moment of force, 4. pressure

OpenStudy (anonymous):

3.moment force